OpenStudy (anonymous):

the maximum height and horizontal range are equal for a projectile launched upward from ground at an angle horizontal equal to

OpenStudy (ujjwal):

this will help here too.. R/H = 4U/V

OpenStudy (anonymous):

options are 60 90 tan^-1/8 tan^-1/4

OpenStudy (anonymous):

plzz help

OpenStudy (ujjwal):

put R=H in above equation U= u cos(theta) V= usin(theta) Find theta..

OpenStudy (anonymous):

i tried but not getting plzzz do this plzz

OpenStudy (ujjwal):

substituting values in above equation \[\theta =\tan^{-1}(4)\] Didn't you get it?

OpenStudy (anonymous):

no plzz can u go it by step by step

OpenStudy (ujjwal):

Put R=H, that makes R/H=1 Now R/H= 4U/V or, 1=4 (u cos(theta)) /(u sin (theta)) that makes, tan (theta) = 4

OpenStudy (anonymous):

hey R=u^2sin2theta/g u have written cos???

OpenStudy (mos1635):

XX' : x=Vo*cosθ*t YY': y=Vo*sinθ*t-1/2*g*t^2 set y=0 and solve to t to find total time of flight put that to x=..... to find x max. YY' Vy=Vo*sinθ-g*t put Vy=0 and solve to t to find time for max height put that t to y=Vo*sinθ*t-1/2*g*t^2 to find h max mast be Xmax=Hmax substitude and see what you will get....

OpenStudy (anonymous):

can some one make it clear

OpenStudy (anonymous):

oh.........no

OpenStudy (anonymous):

can we use R=u^2sin2theta/g H=u^2sin^2theta/2g

OpenStudy (anonymous):

ok

OpenStudy (ujjwal):

Since R=H u^2sin2theta/g =u^2sin^2theta/2g or, 2sin2theta= sin^2 theta or 2*2 sin theta * cos theta = sin^2 theta or, tan theta =4 Do you get it now?

OpenStudy (experimentx):